Circular No. 8476

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005X, 2005Y, 2005Z, 2005aa Further to IAUC 8473, H. G. Khandrika, S. Park, J. Graham, and

W. Li report the LOSS/KAIT discovery of four apparent supernovae on unfiltered KAIT images. SN 2005Z was discovered independently by O. Trondal and M. Schwartz (cf. IAUC 8468) on unfiltered Tenagra II 0.81-m telescope images, and their data are tabulated below.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005X Jan. 24.44 12^{h}24^{m}00.^{s}75 + 7^{o}46′38′′.9 17.5 1′′.6 E, 1′′.8 N
2005Y Jan. 31.12 1 37 37.29 + 0 02 29.8 18.8 2′′.6 E, 5′′.1 N
2005Z Jan. 31.37 10 45 09.18 +22 04 38.3 16.7 5′′.6 W, 4′′.3 S
2005aa Feb. 1.38 9 08 04.21 +27 14 59.4 18.8 10′′.8 W, 0′′.6 S

Additional approximate unfiltered CCD magnitudes, from LOSS/KAIT unless noted otherwise: SN 2005X (in a galaxy that is in turn 5′′.6 east and 27′′.7 south of NGC 4353), 2004 June 10.20 UT, [19.0; Dec. 1.43, [17.5; 2005 Feb. 2.45, 17.7. SN 2005Y in UGC 1159, 2004 Dec. 12.13, [20.0; 2005 Jan. 22.14, [19.0; Feb. 2.23, 18.5. SN 2005Z in NGC 3363, Jan. 15.45, [19.0 (T. Puckett and P. Wiggins); 20.45, [19.0; Feb. 1.46, 17.1; 2.36, 17.3. SN 2005aa in MCG +05-22-8, Jan. 22.39, [19.5; Feb. 2.35, 18.8. Li provides position end figures 09^{s}.07, 37′′.3 for SN 2005Z.

NOTICE TO SUBSCRIBERS AND CONTRIBUTORS The Central Bureau Electronic Telegrams (CBETs) are now available

to subscribers at the CBAT website (http://cfa-www.harvard.edu/iau/

cbet/RecentCBETs.html). While the introductory note on IAUC 8036 indicated the intended ‘temporary’ nature of the CBETs, it was noted on CBET 1 that they may evolve into something more permanent, and discussions with members of the supernova-research community in particular suggest that the CBETs be now expanded into a supplementary publication to these Circulars. One problem with the IAUCs is that they are in a limited-size format due to their printed nature, and this causes problems for long lists of such objects as faint supernovae. The CBETs, being entirely electronic, have no such size constraints, and so they will be henceforth used as a supplementary publication to the IAUCs in that they will often contain data not published on these Circulars; it is anticipated that these Circulars will contain brief announcements of more detailed material appearing in CBETs that is not repeated here, so that there is still at least a printed record of the published topics.

2005 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green